# MetaWFN: A Platform for Unified Implementation of Many-Electron Wavefunctions

### v1

@@ 3 Software Architecture
Redundant codes and repetitive tasks should be avoided as much as possible when developing an extendable and maintainable software package. Therefore, the design philosophy of MetaWFN is to maximize the use of logically identical or similar functionalities. This can be achieved through C++ mechanisms such as polymorphism and template metaprogramming. Specifically, a three-layer abstract architecture (i.e., infrastructure layer, quantum chemistry layer, and algorithm layer; see Fig. 9) is employed in MetaWFN for furnishing a rich set of highly modular and composable functions.
Figure 9: Three-layer architecture of MetaWFN. Each layer abstracts specific components of scientific computation and provides an interface for the layer above. The infrastructure layer abstracts linear algebra operations and parallelization, enabling the design of quantum chemistry and algorithm modules to treat scalar types and parallel strategies as template parameters, thereby handling them uniformly. The quantum chemistry layer encapsulates quantum chemistry information such as Hamiltonian, many-particle basis functions, and molecular symmetry, labeling their properties to facilitate compile-time optimization. Additionally, it provides templated classes to manage molecular integrals, configurations, and Hamiltonian matrix construction. The algorithm layer leverages the quantum chemistry and parallel strategies as template or base classes, describing sophisticated algorithms in MetaWFN in an abstract manner.
@@ 3.1 Infrastructure Layer
The infrastructure layer serves as the basis of MetaWFN. Designed to be independent of quantum chemistry specifics, it provides a unified interface for basic linear algebra operations and algorithm templates for parallel computing, by resolving the dependency of linear algebra libraries on scalar types and the complexity of managing different parallel strategies.
@@ 3.1.1 Linear Algebra Interface
Linear algebra operations are fundamental to computational chemistry algorithms. Optimized libraries like BLAS (Basic Linear Algebra Subprograms) are commonly used for these operations due to their efficiency. However, BLAS functions are explicitly tied to specific scalar types, such as single-precision real (float), double-precision real (double), single-precision complex (std::complex<float>), and double-precision complex numbers (std::complex<double>). Such dependence of scalar types complicates the development of high-level codes that aim to be agnostic of the underlying data types. For example, the standard BLAS function for double-precision matrix multiplication is:
1void cblas_dgemm (const CBLAS_LAYOUT Layout, const CBLAS_TRANSPOSE transa, const CBLAS_TRANSPOSE transb, const MKL_INT m, const MKL_INT n, const MKL_INT k, const double alpha, const double *a, const MKL_INT lda, const double *b, const MKL_INT ldb, const double beta, double *c, const MKL_INT ldc);
In this function, both the function name and parameters are specific to double-precision real numbers. To handle different scalar types, separate functions with different names and parameters are required, leading to code duplication and increased maintenance overhead. To overcome this limitation, MetaWFN introduces a unified linear algebra interface that removes the dependence of scalar types. By utilizing C++ function overloading and templates, we provided a set of functions where the same function name can operate on different scalar types. This allows high-level codes to perform linear algebra operations without concern for the underlying data types. This unification is particularly important because non-relativistic quantum chemical calculations typically involve real numbers, whereas relativistic calculations require complex numbers. By abstracting the scalar types, we laid the groundwork for a Hamiltonian-agnostic implementation, allowing the same high-level code to operate seamlessly on different types of Hamiltonians.
@@ 3.1.2 Parallel Algorithm Template
Efficient parallel computing is crucial for the execution of wavefunction methods. However, managing parallelism, particularly across various strategies such as OpenMP and MPI, can be complex and error-prone. Explicitly addressing these details in each algorithm module results in code duplication and heightens the risk of bugs. To tackle this challenge, MetaWFN utilizes algorithm templates that encapsulate the parallelization logic.
For OpenMP, we provided a base class OpenMPAlgorithmBase that defines several generic parallel execution pattern using OpenMP. This approach allows developers to focus on the algorithmic implementation rather than the intricacies of parallelization, enhancing code cleaning and reducing maintenance overhead. For example, in task-based parallel strategy, the core function of the generic algorithm template is the Run() method, defined as follows:
@@
For the first issue, it is worth noting that OpenMP algorithms rely on built-in dynamic load balancing for loops, but MPI lacks similar functionality. To maximize parallel efficiency, we designed a double-layer dynamic task schedule framework involving both MPI and OpenMP. A unique MPI scheduler on the root process packages for-loop tasks into chunks and dispatches them dynamically to worker nodes. The scheduler monitors worker statuses and assigns the next chunk to the first available node. Each worker unpacks the chunk and processes it using the original OpenMP algorithm. After completing the chunk, the worker signals completion and receives the next chunk. This approach achieves efficient load balancing without algorithm-specific settings and allows the program to adjust dynamically to runtime conditions, such as variations in computation times and hardware differences. The generic OpenMP dynamic schedule is thus divided into two parts in the MPI dynamic schedule, (1) on MPI scheduler:
@@
2MPI_Waitany_process(); // check if any process returns
3MPI_Send_chunk(); // send chunk to idle process
@@
6MPI_Send_stop_sig(); // send stop signal to workers
and (2) on worker nodes:
@@ 2 MPI_Recv()
3if (Recv_calc_sig()) // start another original OpenMP calculation
4{
@@ 5 Run_OpenMP();
6}
7else if (Recv_stop_sig()) // this worker should stop
8{
@@
As for the second issue, we optimized memory usage by utilizing the serialization and memory pool techniques. MPI communication typically requires contiguous memory, but data structures in MetaWFN are automatically generated through inheritance, template combination, and specialization, resulting in complex, non-contiguous memory layouts. To resolve this, we adopted a serialization approach using the cereal library63 for serialization and deserialization. By specifying which class members require serialization, we wrapped MPI functions into template functions that handle serialized streams. This allows our complex data structures to be transmitted over MPI interfaces seamlessly. Consequently, any class with specified serializable members can be used with common MPI functions like MPI_Recv() and MPI_Bcast() automatically.
The use of a memory pool is to control memory usage and facilitate data reuse via pre-allocating large chunks of memory. This technique is inherently compatible with MPI. We categorized memory usage in MetaWFN into static memory and dynamic memory managed by the memory pool. The former is required to store basic information (e.g., configuration spaces) of the program, while the latter refers to intermediate data structures (e.g., ROTs and various records used in the ranking and perturbation steps). In our MPI implementation, the static memory is not distributed; all nodes store a copy. The intermediate data structures in the memory pool are generated from dynamically distributed chunks, which naturally form a distributed memory structure without the need for manual intervention. This approach optimizes memory usage and enhances performance in distributed computing environments.
@@
3.2 Quantum Chemistry Layer
The quantum chemistry layer serves as the core of MetaWFN and is responsible for handling domain-specific components such as Hamiltonian, MPBFs, (abelian) spatial symmetry, (non-abelian) spin symmetry, molecular integrals, and matrix element evaluation. It creates a flexible and extensible framework that can be reused across different quantum chemical methods. A key feature of this layer is the use of a type-trait and tagging system, which allows for the unified handling of various Hamiltonians, molecular symmetries, and MPBFs. By passing information through template parameters and utilizing compile-time mechanisms, both duplication reduction and maintainability enhancement can be achieved. Furthermore, compile-time optimization is also facilitated in this way.
@@ 3.2.1 Type-Trait and Tagging System
To achieve flexibility and extensibility, MetaWFN employs a type-trait and tagging system. This system encodes properties of different Hamiltonians, molecular symmetries, and MPBFs at compile time, allowing the compiler to generate optimized code. As a result, this type-trait and tagging system allows us to write generic code that adapts to different Hamiltonians without code duplication.
A tag is implemented using an empty struct that serve as a unique type representing the specific concept. For example, tags for different Hamiltonians are defined as:
@@
Here, ElectronNonRela and ElectronRela represent the electron-only non-relativistic and relativistic Hamiltonians, respectively, while Hubbard represents the Hubbard model Hamiltonian. These tags are used as template parameters to differentiate between Hamiltonian types in a type-safe manner. To associate properties with these tags, we used type-traits structures (helper classes) that provide compile-time information about the Hamiltonian’s characteristics:
@@
These structures specify whether the Hamiltonian includes certain terms, such as the one-body $H_{1}^{0}+H_{1}^{1}$ and two-body $H_{2}^{0}+H_{2}^{1}+H_{2}^{2}$ terms. To facilitate easy access to these properties, we introduced the following metafunctions [^3] A metafunction is a function or operation that works at compile time rather than at runtime. In C++ programming, especially in template metaprogramming, metafunctions are used to perform computations, checks, or transformations on types and constants at compile time. This allows the compiler to make decisions and optimize code before it runs, which can lead to more efficient and flexible programs.:
@@
These compile-time constants can be used within templates to enable or disable code paths, ensuring that only relevant computations are performed for a given Hamiltonian type. For example, for the Hubbard model, all the code paths related to the evaluation of matrix elements between two-electron excited configurations can be ignored, leading to more efficient implementation. Furthermore, by utilizing the type-trait and tagging system, both spatial and spin symmetries can be readily be handled. As shown in Fig. 11, symmetry adaptation improves significantly the compactness of the variational wavefunctions in selected iCI (SiCI) calculations. Stated differently, on average, a symmetry adapted MPBF tends to contribute more to correlation than the unadapted one (see also Ref. 32).
@@
The type-trait and tagging system also helps to automatically generate a large number of functions with similar computational processes. Consider, e.g., the BCCs $\langle\tilde{I}\mu|e_{\tilde{i}\tilde{j},\tilde{k}\tilde{l}}|\tilde{J}\nu\rangle$ of the reduced CFG pair $(\tilde{I},\tilde{J})$. They involve segment values of all spatial orbitals, under the restriction that outside the interaction region, $[\min(\tilde{i},\tilde{j},\tilde{k},\tilde{l}),\max(\tilde{i},\tilde{j},\tilde{k},\tilde{l})]$, of the operator $e_{\tilde{i}\tilde{j},\tilde{k}\tilde{l}}$, the bra and ket orbitals and their occupations should be the same. The problem lies in that there are seven different types of ROTs and various segment types (cf. Table 2 in Ref. 13). It is extremely tedious and error-prone to write each case individually. This is avoided in MetaWFN by using template functions, where each involved component is determined automatically via the corresponding type-trait and tagging system designed for UGA:
@@
Here, _init_propagate and _final_propagate correspond to the evaluation outside the interaction region, while the other functions handle the evaluation within each region. Specifically, _propagate_segment2 evaluates the orbitals where the operators act, whereas _propagate_segment1 evaluates those where no operator acts. To determine the correct segment value, each function takes TableTy as one of the template parameters and uses the corresponding type-traits for TableTy, along with internal information contained in the class TUGATableMaker. The number specified as the second template parameter distinguishes different regions for a given operator.
Labeling specific concepts via empty structs and extracting/accessing properties via type traiting/metafuncions have been employed in all aspects of MetaWFN, from the lowest infrastructure layer to the highest algorithm layer. This is the essential trick that makes each component independent and the coding as generic as possible.
@@ 3.2.2 Configuration Space
The CfgSpace class in MetaWFN is a central component that manages both the MPBFs and CI vectors. It is determined by 4 template parameters:
@@
The first two parameters are related to CFGs and are hence encapsulated in the BasisMetadata class to represent the MPBFs:
@@ 13 IndxTy indx;
This class stores the essential information of each MPBF, including spin information ($2S$ for CSFs and $2S_{z}$ for DETs), number of open-shell electrons, irreducible representations (irrep), a pointer to the string that stores the CFG, and CSF/DET indices.
The BasisManager class manages a collection of BasisMetadata instances. It provides functionalities for handling MPBFs, such as adding CSFs/DETs, extracting information from a given CSF/DET, or sorting the MPBFs to canonical form. This class is uniform for CSFs and DETs:
@@
Finally, the CfgSpace class represents the entire configuration space, managing both the configurations and the associated CI vectors:
@@
In CfgSpace, the key types, scalar_t, sym_t, indx_t, and others, are determined by the template parameters and propagated via the using declarations. This design ensures that the configuration space is aware of the types used in both the basis manager and vector manager, providing the necessary flexibility for the entire algorithm. The use of template parameters and type binding allows for generic programming, reducing code duplication and enhancing maintainability.
MetaWFN also provides numerous functions, such as printing for debugging purposes, dumping data for restarting computations, merging two configuration spaces, and more. These functionalities enhance the usability and robustness of the MetaWFN package, allowing users to efficiently manage and manipulate the configuration space as needed.

### v2

@@ 3 Software Architecture
Redundant codes and repetitive tasks should be avoided as much as possible when developing an extendable and maintainable software package. Therefore, the design philosophy of MetaWave is to maximize the use of logically identical or similar functionalities. This can be achieved through C++ mechanisms such as polymorphism and template metaprogramming. Specifically, a three-layer abstract architecture (i.e., infrastructure, Hamiltonian, and wavefunction; see Fig. 9) is employed in MetaWave for furnishing a rich set of highly modular and composable functions.
Figure 9: Three-layer architecture of MetaWave. The infrastructure layer abstracts linear algebra operations and parallelization, enabling the Hamiltonian and wavefunction layers to treat scalar types and parallel strategies as template parameters. The Hamiltonian layer encapsulates all information relevant to the Hamiltonian matrix, including molecular symmetry, MO integrals, MPBFs, and matrix element evaluation. The wavefunction layer leverages the Hamiltonian and infrastructure layers for unified implementation of relativistic and nonrelativistic many-electron wavefunctions.
@@ 3.1 Infrastructure Layer
The infrastructure layer serves as the basis of MetaWave. Designed to be independent of quantum chemistry specifics, it provides a unified interface for basic linear algebra operations and algorithm templates for parallel computing, by resolving the dependency of linear algebra libraries on scalar types and the complexity of managing different parallel strategies.
@@ 3.1.1 Linear Algebra Interface
Linear algebra operations are fundamental to computational chemistry algorithms. Optimized libraries like BLAS (Basic Linear Algebra Subprograms) are commonly used for these operations due to their efficiency. However, BLAS functions are explicitly tied to specific scalar types, such as single-precision real (float), double-precision real (double), single-precision complex (std::complex<float>), and double-precision complex numbers (std::complex<double>). Such dependence of scalar types complicates the development of high-level codes that aim to be agnostic of the underlying data types. Take the standard BLAS function for double-precision matrix multiplication (Listing 1) as an example. In cblas_dgemm both the function name and parameters are specific to double-precision real numbers. To handle different scalar types, separate functions with different names and parameters are required, leading to code duplication and increased maintenance overhead. To overcome this limitation, MetaWave introduces a unified linear algebra interface that removes the dependence of scalar types. By utilizing C++ function overloading and templates, we provided a set of functions where the same function name can operate on different scalar types. This allows high-level codes to perform linear algebra operations without concern for the underlying data types.
Listing 1: Declaration of cblas_dgemm.
1void cblas_dgemm(const CBLAS_LAYOUT Layout,
2 const CBLAS_TRANSPOSE transa,
3 const CBLAS_TRANSPOSE transb,
4 const MKL_INT m, const MKL_INT n, const MKL_INT k,
5 const double alpha, const double *a, const MKL_INT lda,
6 const double *b, const MKL_INT ldb, const double beta,
7 double *c, const MKL_INT ldc);
This unification is particularly important because nonrelativistic quantum chemical calculations typically involve real numbers, whereas relativistic calculations require complex numbers. By abstracting the scalar types, we laid the groundwork for a Hamiltonian-agnostic implementation, allowing the same high-level code to operate seamlessly on different types of Hamiltonians.
@@ 3.1.2 Parallel Algorithm Template
Efficient parallel computing is crucial for the execution of wavefunction methods. However, managing parallelism, particularly across various strategies such as OpenMP and MPI, can be complex and error-prone. Explicitly addressing these details in each algorithm module results in code duplication and heightens the risk of bugs. To tackle this challenge, MetaWave utilizes algorithm templates that encapsulate the parallelization logic.
For OpenMP, we provided a base class OpenMPAlgorithmBase that defines several generic parallel execution pattern using OpenMP. This approach allows developers to focus on the algorithmic implementation rather than the intricacies of parallelization, enhancing code cleaning and reducing maintenance overhead. For example, in task-based parallel strategy, the core function of the generic algorithm template is the Run() method (Listing 2).
Listing 2: OpenMP parallel algorithm template in MetaWave.
@@
For the first issue, it is worth noting that OpenMP algorithms rely on built-in dynamic load balancing for loops, but MPI lacks similar functionality. To maximize parallel efficiency, we designed a double-layer dynamic task schedule framework involving both MPI and OpenMP. A unique MPI scheduler on the root process packages for-loop tasks into chunks and dispatches them dynamically to worker nodes. The scheduler monitors worker statuses and assigns the next chunk to the first available node. Each worker unpacks the chunk and processes it using the original OpenMP algorithm. After completing the chunk, the worker signals completion and receives the next chunk. This approach achieves efficient load balancing without algorithm-specific settings and allows the program to adjust dynamically to runtime conditions, such as variations in computation times and hardware differences. Thus, the generic MPI dynamic schedule is devided into two parts, the MPI scheduler (Listing 3) and work nodes (Listing 4).
Listing 3: Scheduler-side code for MPI dynamic scheduling in MetaWave.
@@
2 MPI_Waitany_process(); // check if any process returns
3 MPI_Send_chunk(); // send chunk to idle process
@@
6MPI_Send_stop_signal(); // send stop signal to workers
Listing 4: Work nodes code for MPI dynamic scheduling in MetaWave.
@@ 2 MPI_Recv()
3 if (Recv_calc_signal()) // start another original OpenMP calculation
4 {
@@ 5 Run_OpenMP();
6 }
7 else if (Recv_stop_signal()) // this worker should stop
8 {
@@
As for the second issue, we optimized memory usage by utilizing the serialization and memory pool techniques. MPI communication typically requires contiguous memory, but data structures in MetaWave are automatically generated through inheritance, template combination, and specialization, resulting in complex, non-contiguous memory layouts. To resolve this, we adopted a serialization approach using the cereal library72 for serialization and deserialization. By specifying which class members require serialization, we wrapped MPI functions into template functions that handle serialized streams. This allows our complex data structures to be transmitted over MPI interfaces seamlessly. Consequently, any class with specified serializable members can be used with common MPI functions like MPI_Recv() and MPI_Bcast() automatically.
The use of a memory pool is to control memory usage and facilitate data reuse via pre-allocating large chunks of memory. This technique is inherently compatible with MPI. We categorized memory usage in MetaWave into static memory and dynamic memory managed by the memory pool. The former is required to store basic information (e.g., configuration spaces) of the program, while the latter refers to intermediate data structures (e.g., ROTs and various records used in the ranking and perturbation steps). In our MPI implementation, the static memory is not distributed; all nodes store a copy. The intermediate data structures in the memory pool are generated from dynamically distributed chunks, which naturally form a distributed memory structure without the need for manual intervention. This approach optimizes memory usage and enhances performance in distributed computing environments.
@@
3.2 Hamiltonian Layer
The Hamiltonian layer serves as the core of MetaWave and is responsible for handling domain-specific components such as Hamiltonian, MPBFs, (abelian) spatial symmetry, (non-abelian) spin symmetry, molecular integrals, and matrix element evaluation. It creates a flexible and extensible framework that can be reused across different quantum chemical methods. A key feature of this layer is the use of a type-trait and tagging system, which allows for the unified handling of various Hamiltonians, molecular symmetries, and MPBFs. By passing information through template parameters and utilizing compile-time mechanisms, both duplication reduction and maintainability enhancement can be achieved. Furthermore, compile-time optimization is also facilitated in this way.
@@ 3.2.1 Type-Trait and Tagging System
To achieve flexibility and extensibility, MetaWave employs a type-trait and tagging system. This system encodes properties of different Hamiltonians, molecular symmetries, and MPBFs at compile time, allowing the compiler to generate optimized code. As a result, this type-trait and tagging system allows us to write generic code that adapts to different Hamiltonians without code duplication, in the same spirit as other codes45, 78, 79, 80.
A tag is implemented using an empty struct that serve as a unique type representing the specific concept. For example, tags for different Hamiltonians are defined in Listing 5, where ElectronNonRela and ElectronRela represent the electron-only nonrelativistic and relativistic Hamiltonians, respectively, while Hubbard represents the Hubbard model Hamiltonian. These tags are used as template parameters to differentiate between Hamiltonian types in a type-safe manner.
Listing 5: Hamiltonian tag system in MetaWave.
@@
To associate properties with these tags, we used type-traits structures (helper classes) that provide compile-time information about the Hamiltonian’s characteristics (Listing 6).
Listing 6: Hamiltonian type-traits system in MetaWave.
@@
These structures specify whether the Hamiltonian includes certain terms, such as the one-body $H_{1}^{0}+H_{1}^{1}$ and two-body $H_{2}^{0}+H_{2}^{1}+H_{2}^{2}$ terms. To facilitate easy access to these properties, we introduced the metafunctions [^3] A metafunction is a function or operation that works at compile time rather than at runtime. In C++ programming, especially in template metaprogramming, metafunctions are used to perform computations, checks, or transformations on types and constants at compile time. This allows the compiler to make decisions and optimize code before it runs, which can lead to more efficient and flexible programs. (Listing 7).
Listing 7: Hamiltonian metafunctions in MetaWave.
@@
These compile-time constants can be used within templates to enable or disenable code paths, ensuring that only relevant computations are performed for a given Hamiltonian type. For example, for the Hubbard model, all the code paths related to the evaluation of matrix elements between doubly excited configurations can be ignored, leading to a more efficient implementation. Moreover, by utilizing the type-trait and tagging system, both spatial and spin symmetries can be readily be handled. As shown in Fig. 11, symmetry adaptation improves significantly the compactness of the variational wavefunctions in selected iCI (SiCI) calculations. Stated differently, on average, a symmetry adapted MPBF tends to contribute more to correlation than the unadapted one (see also Ref. 35).
@@
The type-trait and tagging system also helps to automatically generate a large number of functions with similar computational processes. Consider, e.g., the BCCs $\langle\tilde{I}\mu|e_{\tilde{i}\tilde{j},\tilde{k}\tilde{l}}|\tilde{J}\nu\rangle$ of the reduced oCFG pair $(\tilde{I},\tilde{J})$. They involve segment values of all spatial orbitals, under the restriction that outside the interaction region, $[\min(\tilde{i},\tilde{j},\tilde{k},\tilde{l}),\max(\tilde{i},\tilde{j},\tilde{k},\tilde{l})]$, of the operator $e_{\tilde{i}\tilde{j},\tilde{k}\tilde{l}}$, the bra and ket orbitals and their occupations should be the same. The problem lies in that there are seven different types of ROTs and various segment types (cf. Table 2 in Ref. 16). It is extremely tedious and error-prone to write each case individually. This is avoided in MetaWave by using template functions, where each involved component is determined automatically via the corresponding type-trait and tagging system designed for UGA.
Take the function (Listing 8) for generating the BCCs between doubly excited oCFG pair $(\tilde{I},\tilde{J})$ as an example. Given the input parameter csfindx (which corresponds to $\nu$), it generates all non-vanishing BCCs $\langle\tilde{I}\mu|e_{\tilde{i}\tilde{j},\tilde{k}\tilde{l}}|\tilde{J}\nu\rangle$ and store the results and the indices $\mu$ in drct and indx respectively. If exists, the function also generates $\langle\tilde{I}\mu|e_{\tilde{i}\tilde{l},\tilde{k}\tilde{j}}|\tilde{J}\nu\rangle$ and stores it in xchng Since the number of CSFs that one oCFG can generate cannot exceed the range of int32_t in routine applications, int32_t is used to represent the indices $\mu$ and $\nu$. In maketable_dbl, _init_propagate and _final_propagate correspond to the evaluation outside the interaction region, while the other functions handle the evaluation within each region. Specifically, _propagate_segment2 evaluates the orbitals where the operators act, whereas _propagate_segment1 evaluates those where no operator acts. To determine the correct segment value, each function takes TableTy as one of the template parameters and uses the corresponding type-traits for TableTy, along with internal information contained in the class TUGATableMaker. The number specified as the second template parameter distinguishes different regions for a given operator.
Listing 8: Template function for generating BCCs between doubly excited oCFG pairs.
@@
Labeling specific concepts via empty structs and extracting/accessing properties via type traiting/metafuncions have been employed in all aspects of MetaWave, from the lowest infrastructure layer to the highest wavefunction layer. This is the essential trick that makes each component independent and the coding as generic as possible.
@@ 3.2.2 Configuration Space
The CfgSpace class in MetaWave is a central component that manages both the MPBFs and CI vectors. It is determined by 4 template parameters:
@@
The first two parameters are related to oCFGs and are hence encapsulated in the BasisMetadata class (Listing 9) to represent the MPBFs. This class stores the essential information of each MPBF, including spin information ($2S$ for CSFs and $2S_{z}$ for DETs), number of open-shell electrons, irreducible representations (irrep), a pointer to the string that stores the oCFG, and CSF/DET indices.
Listing 9: Definition of class BasisMetadata representing the MPBFs.
@@ 13 IndxTy indx;
The BasisManager class (Listing 10) manages a collection of BasisMetadata instances. It provides functionalities for handling MPBFs, such as adding CSFs/DETs, extracting information from a given CSF/DET, or sorting the MPBFs to canonical form. This class is uniform for CSFs and DETs.
Listing 10: Definition of class BasisManager.
@@
Finally, the CfgSpace class (Listing 11) represents the entire configuration space, managing both the configurations and the associated CI vectors. In CfgSpace, the key types, scalar_t, sym_t, indx_t, and others, are determined by the template parameters and propagated via the using declarations. This design ensures that the configuration space is aware of the types used in both the basis manager and vector manager, providing the necessary flexibility for the entire algorithm. The use of template parameters and type binding allows for generic programming, reducing code duplication and enhancing maintainability.
Listing 11: Definition of class CfgSpace.
@@
MetaWave also provides numerous functions, such as printing for debugging purposes, dumping data for restarting computations, merging two configuration spaces, and more. These functionalities enhance the usability and robustness of the MetaWave package, allowing users to efficiently manage and manipulate the configuration space as needed.